A company's board of directors wants to form a committee of
2 of its members. There are 7 members to choose from. How many different committees of 2 members could possibly be formed?

Answer by fractalier(6550)

7 chosen 2 at a time is found by
7! / (7-2)!(2!) = 7*6 / 2*1 = 21 ways